OpenStudy (anonymous):

Simplify the expression. Use positive exponents. 1. m^3n–6p^0

Nnesha (nnesha):

exponent rules ! everything to the 0 power is just = 1 \[\large\rm x^0 = 1\]\[\large\rm (459u504xxx^4 23)^0 = 1\] \[\large\rm (ab^5cxyz^6)^0 = 1\]

Nnesha (nnesha):

exponent rule! \[\huge\rm x^{-m} = \frac{ 1 }{ x^m }\] if there is negative exponent change that to its reciprocal

OpenStudy (anonymous):

so the answer is m^3/n^6

Nnesha (nnesha):

:-)

Nnesha (nnesha):

:-) = that's right

Nnesha (nnesha):

btw \(\huge\color{Green}{{\rm welcome}\rm~to~open~study!!!!}\)\(\Huge \color{gold}{\star^{ \star^{\star:)}}}\Huge \color{green}{\star^{ \star^{\star:)}}}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\)\(\rm\color{green}{o^\wedge\_^\wedge o}\)

OpenStudy (anonymous):

Thank you ^_^

Nnesha (nnesha):

my pleasure :-)